Circular No. 7341

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000A IN MCG +1-59-81 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126), reports the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a supernova of mag ∼ 16.5 on unfiltered images taken on Jan. 1.07 and 1.20 UT, confirmed at about the same brightness on Jan. 2.10. SN 2000A is located at α = 23^{h}27^{m}58^{s}.90, δ = +8^{o}47′02′′.1 (equinox 2000.0), which is 1′′.6 east and 4′′.6 north of the nucleus of MCG +1-59-81, which in turn is 30′′ northeast of NGC 7674. A KAIT image of the same field taken on 1999 Dec. 22.1 shows nothing at the position of SN 2000A (limiting mag ∼ 19.0).

S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000A, taken by P. Berlind on Jan. 3.08 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-Ia supernova at maximum light. Narrow superimposed Hα emission gives a recession velocity for the host galaxy (MCG +1-59-81) of 8760 km/s, which matches the NASA/IPAC Extragalactic Database recession velocity for NGC 7674, indicating that the supernova’s host galaxy and NGC 7674 are associated. The supernova photospheric expansion velocity, measured to the Si II (rest 635.5 nm) absorption trough, is 9500 km/s.

SUPERNOVA 1999gp IN UGC 1993 S. Jha et al. also report that a spectrum of SN 1999gp (cf. IAUC 7337),

obtained by P. Berlind on Jan. 2.21 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum light. The absence of a strong Si II 580-nm feature may indicate that the supernova is overluminous. Narrow superimposed lines of Hα yield a recession velocity of 8300 km/s for the host galaxy. The supernova expansion velocity, measured to the Si II (rest 635.5 nm) absorption trough, is 11 500 km/s.

SUPERNOVA 1999gh IN NGC 2986 Unfiltered CCD magnitudes by S. Yoshida and K. Kadota, Ageo,

Saitama, Japan (comparison-star magnitudes from Tycho V data): 1999 Dec. 6.775 UT, 14.7; 9.781, 15.0; 27.765, 15.6.

2000 January 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green